IT for FIT Servicing > Current Downtime Schedule
REMINDER: The 1st and 3rd Thursday of every month from 7am to 9am our applications servers will
receive updates/patches and preventive maintenance.
For most users, these are www.fitnyc.edu, webreg.fitnyc.edu, imail.fitnyc.edu, fitdil.fitnyc.edu, MyFIT.fitnyc.edu and the library.
No more than one third of our servers will be affected at any one time. In the future you will recieve updates on the MyFIT portal and by email.
Thursday, December 11, 2008
From 7:30am to 11am, Banner and Banner student Information
system will be unavailable due to mandatory software and
Please note that student and employee access to Banner
information via WebReg or the MyFIT portal will be closed
and the Online Information System, timesheets, class
rosters, Xtender, schedules, NolijWeb, Housing Web, Brio
reports, online bill paying, registration and Banner INB
will all be unavailable.
We are sorry for any inconvenience this may cause.
Top of Page
Thursday, November 6, 2008
Banner, Webreg, Online Information System, Bitech, Xtender systems are unavailable. Construction work in the data center disrupted storage systems.
Firmware upgrade of Banner/Bitech/Bantest storage array;
Webreg Switchover to new server; MS Patch group 3 updates; 6
Ironport/Baracuda Firmware Updates; Financial Aid Patches
applied to Banner Prod; General Counsel Xtender Application
into production and Apply dual Luminus roles for precollege
Wednesday, November 5, 2008
FIT is currently experiencing delays in our outbound mail
service. We are currently working with the hardware vendor
as to repair and restore services.
Tuesday, November 4, 2008
Due to the renovation work being done in C100 one of the
rj45 cables got dislodged. Access to the internet to be
interrupted for a short period.
Top of Page
Thursday, October 30, 2008
Replacement of the 7206 Edge Router 6:00 am – 7:00 am
Removal and replacement of the 7 (DMZ) switch - 7:00 am – 8:00
MS Patch group 2 updates - 8:00 am - 9:00 am
Thursday, October 23, 2008
Firmware upgrade of Luminis storage array
October 16, 2008
Replacement of router parts in 236 during the hours of 7am
Tuesday, October 07, 2008
Bi-Tech taken down for residual maintenance this morning
between the hours of 10: 00 -11:00.
Monday, October 06, 2008
Due to a hardware failure Bi-Tech and Banner are not
Thursday-Friday, October 2 - 3, 2008
On Tuesday night after the C108 electrical shutdown, the
Bi-Tech database became corrupted and data entered into the
system on Monday 9/29 from 4:30 to 6:00 PM was lost. Any
work completed in Bi-Tech during this time will need to be
Friday, October 3, 2008
Webfiles.fitnyc.edu and Faculty/Student "N" drives
Top of Page
Friday, September 19, 2008
Late Thursday evening Xythos Tomcat server on 220.127.116.11
not available - services were restored at 8:00am.
Thursday, September 18, 2008
Due to emergency system maintenance Banner was not available
8am - 11am.
Tuesday, September 16, 2008
Students were not able to save their work on Xythos after
several attempts the Xythos drives were brought up on both
app-v445-26 And app-v445-27.
During the construction of the A labs project, there were 12
data lines accidentally cut. The areas affected were the
Academic skills department and the International Student
Advisement. Gilbane, the company on site, will re-run the
cables tomorrow morning at 7am. Telecom will terminate the
cables in the closet after they are run
Saturday, September 13, 2008
The UPS crashed on Saturday, the power supply on the BCMS PC
server burnt out. The power supply was replaced on Tuesday
Thursday, September 11, 2008
Unscheduled: The S and T drives (transport) were not available at 1:53 PM they were available for Faculty and Students 10:28 PM.
Scheduled: IT for FIT is in the process of renovating the college’s data center. As a result, many technology resources in all academic and residential life buildings on West 27 Street will be unavailable from 7:00 pm on Saturday, September 13 through approximately 7:00 am on Sunday, September 14. Internet, email, Banner, Bi-Tech, and telephones will not be operational. Cable television in the dormitories will not be affected.
Building lobbies will have working telephones in the event of an emergency.
During the hours of 8 AM - 9AM the following patches and
upgrades will be done - we do not expect any interruptions.
MS patch Group 2; Synchronization of @ numbers in Xtender
with Spriden records.; Imail removal of high school student
Thursday, September 04, 2008
Scheduled: During the hours of 8:00 am - 10:00 am Block
Section upgrades installed in Banner, New applicant profile
instaslled in WebReg and Luminis.
Unscheduled: WebFiles/ Xythos not available 10:00 PM -
1:00AM, the Java processes used to control Xythos stopped
functioning - manually generated a core file and restarted
the Tomcat/Java service.
Tuesday, September 02, 2008
Monday, September 01, 2008
Webfiles/Xythos unavailable at 6pm
MyFIT experienced channel errors
Top of Page
Friday, August 29, 2008
Exchange unable to send mail 9:00am - 11:20am.
Brio is currently up but experiencing technical problems. IT
is working on the issue.
Monday August 25, 2008
Students in the 31st Street Residence Hall do not have
access to wireless and wired connections to the internet.
Thursday August 21, 2008
Wednesday August 20, 2008
From 8am – 10:30am email and Outlook was not available and
from 8am – 11am due to the mailbox cluster being down MyFIT
portal was not available 11:00am – 1:00 due to emergency
From 8-10am the telephone system had to have the T1 line reprogrammed
during this time no phone service was available.
Thursday, August 14, 2008
Top of Page
The Xythos Oracle databases was taken down to configure a
higher number of maximum database connections. Therefore,
unavailable until the databases were brought back on line.
Thursday, July 24, 2008
Patch Group 3 updates
31st Residence Hall network software upgraded during the
hours of 9:30 am – 10:30 am; during this time the 31st
Street students might experience intermittent interruptions
when connecting to the internet.
Friday, July 18, 2008
College telephone system – unit
will be replaced between 6:00 pm - 9:00 pm. We do not anticipate
that there will be any interruption in telephone service.
However, there is a possibility that some users may
experience brief or intermittent delays making connections.
Thursday, July 17, 2008
Bi-Tech upgrade 8:00am - 11:00am during
this time Bi-Tech will not be available. The MERU controller
will be upgraded Thursday July 17th and July 18th during the
hours of 8:00am – 11:00am there will be no wireless access.
Wednesday, July 16, 2008
Sunday July 13, 2008
There was a significant delay in Mail
processing on the IMAIL server which also affected exchange
users. As of this email IMAIL is now processing email and
the inbound and outbound queues are being processed.
Week of June 30 - Email problems due to barracuda
configuration and license issue
Thursday, July 10, 2008
Patch Group 2 updates - Xythos Solaris patch updates to the
July 10th list
Tuesday, July 01, 2008Top of Page
8:00am -1:00pm Bi-Tech accounting updates system not
available from 8am -1pm.
- Baracuda SMTP firmware updated
Thursday, June 26, 2008
Patch Group I updates
Thursday, June 19, 2008
Baracuda SMTP firmware update
Residence Halls on 27th street will not have internet
services between the hours of 8am – 12pm due to the Bradford
Monday, June 16, 2008
Banner will not be available from 8 AM – 12 PM. Please note that the Online
Information System, timesheets, class rosters, Xtender,
schedules, Nolij, Housing Web, Brio reports, online bill
paying, student and employee access to Banner information
via the MyFIT Portal and registration will all be
The old link to Banner INB on your desktop will no
longer function. After the upgrade, to access Banner INB,
you MUST go to
login, and then choose the employee tab and your access to
Banner INB will be via a link listed there. Banner INB will
now be available only through the MyFIT Portal will continue
only to be available from on campus.
Saturday, June 14, 2008
Due to a hardware failure, staff and administrative "N"
drives are unavailable. Additionally, some departmental
shares or “M” drives are also affected. A restore is
underway, but due to the volume of data, the process
requires that the system remain unavailable to all staff and
administrative employees during this time. IT expects to
have the restore completed later this week. Please note that
the faculty and student “N” drives are available.
Thursday, June 12, 2008
31st street residence hall will not have internet services
between the hour of 8am -12pm due to the controller being
Thursday, June 5, 2008
The following systems will be updated; Baracuda SMTP firmware, Ironport SMTP
firmaware, Patch Group, Financial aid updates, and BiTech Health Insurance.
Reconfigure 27th to 31st VPN 11am to 2pm
Wednesday, June 4, 2008
At 10am - Due to a catastrophic electrical failure, the Xythos Application is
temporarily unavailable. We are diligently working on correcting/bypassing the
Tuesday, June 3, 2008
9:30 am, there was a disruption with Campus Wide Printing which necessitated a
system restart. Printing Services have been restored.
Top of Page
Friday, May 23, 2008
The Bitech Prod Server (18.104.22.168) was
rebooted at 8:00 pm that night.
Thursday, May 22, 2008
7am – 8am Patch Group
2; UPS Maintenance; SUNYPerson & EDUPerson SUNY required LDAP schema change VPN
move to 9 segment – due to integration issues, adv coordination With SUNYNet
3:00 PM the telephones in C100 and
Admissions went down. Troubleshooting revealed the 6506 in
C100 to have partially shut down due to overheating. Further
investigation determined the mounting method to be
responsible for the shutdown. No air at all was able to move
thru the system, the system was not able to cool itself. The
6506 from the rack was removed and used a saws all to change
the mounting configuration resulting in the establishment
air flow thru the system. Telephone service was restored at
approximately 3:45 PM. This solution is temporary!
The air flow is still inhibited. The correct mounting
bracket must be obtained and properly installed.
Saturday, May 17, 2008
9:45AM - the Saflok card reader could not connect to the server; the PC was reconfigured
to connect to the server. Services were available at 11:30AM.
Friday, May 16, 2008
The ancient Cisco 2900 servicing room C-306 failed completely, shortly after
lunch today. It has been replaced with a 24 port Cisco 3550 and tested fine. The
service outage was almost 1 hour.
Thursday, May 15, 2008
No downtime scheduled due to finals and final projects. Off
campus access to the exchange email platform was unavailable for 25 minutes due
to emergency maintenance.
Tuesday, May 06, 2008
Due to the planned maintenance of an air conditioning unit inside C-108, the
ambient air temp inside the computer center peaked at about 8:47 a.m..
This caused a protection fault on a 48 port ethernet blade servicing the users
in C-100 and perhaps a few others in C building (first floor). The blade powered
down and cooled and was reset with a few minutes restoring service to those
users. Since this all occurred before 9 a.m., many were unaware of these events.
We've been advised that the temperature issues have been resolved and we don't
expect this issue to reoccur.
Monday, May 05, 2008
The ADMITMAC server has encountered an Operating System failure. ETS is working
with ICC in addressing this issue .
Wireless users on the south side of the campus are having problems getting IP
addresses. In order to restore service, the Bradford Campus Manager has been
bypassed for this area, which also acts as the DHCP server for this group. Users
are getting IP addresses directly from the enterprise DHCP server.
Thursday, May 01, 2008
This afternoon, Pharos Pay-For-Print services was hung for about 10 minutes this
afternoon. The problem occurred at 2:42PM this afternoon and cleared out
automatically by 2:52PM.
Top of Page
Wednesday, April 30, 2008
Brio Production is down until later this afternoon due to system hardware
issues. We will keep you advised of the status. If you have an emergency
situation where you must run a Brio production report or process please contact
Judy Duncan-Allen in IT. The raid card was replaced and the system is now up and
running from 4/30/08 12:24 pm
Tuesday, April 29th, 2008
appsc10804 / brio down - the raid card was
replaced and the system is now up and running from 4/30/08 12:24 pm
Thursday, April 24, 2008
From 7:00 am – 8:00 am - Cutover Meru 110 controller to 3150 this will impact
the south side of 27th street. During this time there will be no wireless access
available in the Museum, Res Life Alumni work Room, Conference Room, the Res
Life Co-ED lounge/rec room and 236 W. 27th St. on the 1st and 2nd floor.
The cutover was not successful and the reconfiguration of the controllers are
still in progress as of 4/30/08
Friday, April 18th, 2008
Blackboard will not be available on Friday, April 18th from 2am until 6am,
during this time the system will be upgraded.
Thursday, April 17th, 2008
The following systems will not be available
during the hours of 8:00am – 10pm; Ironport will have patches and security
updates applied as will Patch Group 3.
Wednesday, April 16th, 2008
The Bi-Tech upgrade has been successfully deployed and the system is now available.
There are new features and fixes available in the upgraded version of Bi-Tech:
1.Link to attachments in the approval email (in the Attached Documents section; the bottom of the email); The approver will still need to login to Bi-Tech, but it is not necessary to go to the Task list to view attachments.
2.Ability to see future workflow tasks (the future tasks will appear as TBA till they will reach the approvers queue).
3.Budget transfer object code issue has been resolved.
Unfortunately there is also a bug that we discovered:
When the Requisition/Voucher is resubmitted to workflow using C1, the creator’s name is showing up under the approver’s name. Once the Requisition/Voucher goes through the approval process creator’s name will be removed from the list of approvers. You will not need to re-approve the requisition/voucher and it will go through the appropriate approval process.
We reported this issue to SunGard Bi-Tech and we are working with them to have it resolved.
Monday through *Wednesday*
Thursday, April 14-17, 2008
Bi-Tech system will not be available from 8am on Monday, April 14th till 12pm on
Thursday, April 17th. The system will be off-line for this limited time to allow
for the system upgrade.
Please complete any time-sensitive transactions in Bi-Tech prior to this time.
Top of Page
Thursday, March 27, 2008
On Thursday, March 27, from 4pm to 5pm, Bi-Tech will not be
available, the system will be off-line for this limited time
to allow for a system upgrade. Please complete any
time-sensitive transactions in Bi-Tech prior to that time.
Tuesday, March 18, 2008
On Tuesday, March 18, 2008, IT will implement a mandatory
upgrade of the Banner Oracle Database. To complete this
upgrade, Banner will be unavailable all day. Please schedule
any business activities requiring access to Banner
accordingly. This technology upgrade is a requirement from
SunGard Higher Education and must be completed by April 2008
for the College to be prepared to apply the mandated summer
Banner Financial Aid upgrade.
Also from 8am to noon, BiTech will be unavailable due to
system updates. All servers in “Group1” are patched with
Microsoft Security Updates except for FITMAIL and
TB-C108-01. A backup involving these two servers are
currently in progress.
Monday, March 17, 2008
On Monday, March 17, there was a MS Exchange, SP1
installation, that lasted for 1 hour.
Thursday, March 6, 2008
On Thursday, March 6, from 8am to noon, Banner will be
unavailabe due to software upgrades. Please note that the
Online Information System, timesheets, class rosters,
Xtender, schedules, Nolij, housing web, Brio reports, online
bill paying, registration and Banner INB will not be
available. In addition, student and employee access to
Banner information via the MyFIT portal will also be
Top of Page
Thursday, February 21, 2008
On Thursday, February 21, from 6pm to 8pm, Bi-Tech system
will not be available. The system will be off-line for this
limited time to allow for the upgrade. Please complete any
time-sensitive transactions in Bi-Tech prior to that time.
February 7 – February 20, 2008
No scheduled downtime – due to first week of classes
Friday, January 18, 2008
On Friday, January 18, Student File (N drives), Class Drop
Box Purge files were emptied for Spring 08 semester.
Thursday, January 17, 2008
On Thursday, January 17, patched group 1: PassagePoint,
Backup_Server, Domain_Controller, WorkGroupShare,
File_Server Staff, NFS Shares/'U' Drive, Old_Payment System,
WebPDMApps, WebPDM, JDA/ARTHUR, BANNER X DRIVE, Paper
Vision/Digitech system, Mail Server and Students' 'N' Drive.
Monday, January 14, 2008
On Monday, January 14, reset Bradford database.
Friday, January 4, 2008
On Friday, January 4, from 9:00 am through 1:00 pm, DARS the Degree Audit Reporting System will be unavailable due to updates to the system.
Top of Page